OpenStudy (anonymous):

Solve the system of equations. 5x - y = 3 and -10x + 2y = -6

OpenStudy (anonymous):

do you prefer substitution or elimination ?

OpenStudy (wwe123):

-y =-5x+3 y=5x-3 eq2 -10x +2(5x-3)=-6 -10x+10x-6 =-6 =0

OpenStudy (anonymous):

substitution is probably easier.....let me show you... 5x - y = 3 (isolate y by subtracting 5x from both sides) -y = -5x + 3 (multiply by -1 to make y positive) y = 5x - 3 now sub 5x - 3 in for y in the other equation and solve for x. Once x is found, sub that value back into either equation and solve for the numerical value of y.

OpenStudy (anonymous):

The answer choices (0, 5) (-1, 8) (2, -1) no solution infinitely many solutions

OpenStudy (anonymous):

if your result comes out 0 = 0 or 1 = 1, or something equal like that, then there is infinite solutions. If it comes out like 1 = 2 or 2 = 3 or something not equal, then there is no solution. And of course, if you get x = a number and y = a number, then that set is the answer (one solution)

OpenStudy (anonymous):

wait, so whats the answer ? @kelliegirl33

OpenStudy (anonymous):

infinitely many solutions